A positive particle of charge 1.0C accelerates in a uniform electric field of 100 V/m. The particle started from rest on an equipotential plane of 50V. After t=0.0002 seconds, the particle is on an equipotential plane of V=10 volts. Determine the distance travelled by the particle.
1
Expert's answer
2020-12-17T09:11:16-0500

1) "W=\Delta KE=-q\Delta \phi=-1\cdot(10-50)=40(J)"


"W=Fd=qEd"


"qEd=40\to d=40/(qE)=40/(1\cdot100)=0.4(m)"


or


2) "E=-\frac{\Delta\phi}{d}\to d=-\frac{\Delta\phi}{E}=-\frac{10-50}{100}=0.4(m)"